# https://perso.usthb.dz/~slarabi/accueil.html

### v1

@@ 1 Introduction
The remainder of this paper is organized as follows: Section 2 discusses related works in the domain of machine thinking. Section 4 describes the proposed framework, detailing its components and functionalities. Section 6 presents experimental results, and finally we conclude in section  7 this paper.
@@ 5.2 Thinking: A Computational Perspective
Once a sentence (representing an event) is communicated by the CTU unit, a set of mental images is generated and visualized, representing the progression from the initial state to the final state. These visualized images correspond to the unfolding of the event from beginning to end. The generation and visualization of these images rely on contextual knowledge, inference, and the ability to introduce changes to the content of the images. The change of the mental image content is made in order to add details, or to correct partially the event or change completely the content. Once the process of seeing and adjusting the content of mental images, the event is scheduled as a set of elementary events corresponding to the set of generated mental images.
@@ 6.1 Input Data: available and generated
The Needs Unit contains the actions to be performed expressed by the CTU unit and scheduled.
@@ 6.3 Draw the Mental Images
Reasoning can be applied to each image by asking logical questions—for example, ’What if the key doesn’t open the door?’ The CTU unit proposes a new action:"the person is nervous" , then a new mental image is then generated (see figure 8). The same process is repeated and a new possible actions are generated: "the person breaks down the door", "the person call the firefighters for they open the door". Mental images generated are shown by figure 8.
@@ 7 Conclusion
In this paper, we addressed the ambitious challenge of enabling machine thinking by proposing a comprehensive framework inspired by recent advancements in neuroscience, deep learning, and artificial intelligence. While significant progress has been made in developing systems capable of processing sensory data, reasoning, and interacting with humans, our approach extends beyond existing models by integrating the concept of an Mental Image to simulate machine cognition.

### v2

@@ 1 Introduction
The remainder of this paper is organized as follows: Sections 2 and 3 discuss related work in the domains of machine thinking and mental imagery. Sections 4 and 5 describe the proposed framework, detailing its components, functionalities, and how it performs reasoning through the generation of mental images. Section 6 presents the experimental results, and finally, Section 7 concludes the paper.
@@ 5.2 Thinking: A Computational Perspective
Once a sentence (representing an event) is communicated by the CTU unit, a set of mental images is generated and visualized, representing the progression from the initial state to the final state. These visualized images correspond to the unfolding of the event from beginning to end (see figure 4). The generation and visualization of these images rely on contextual knowledge, inference, and the ability to introduce changes to the content of the images. The change of the mental image content is made in order to add details, or to correct partially the event or change completely the content. Once the process of seeing and adjusting the content of mental images is complete, the event is scheduled as a set of elementary events corresponding to the set of generated mental images.
@@ 6.1 Input Data: available and generated
The Needs Unit contains the scheduled actions to be performed by the CTU unit.
@@ 6.3 Draw the Mental Images
Reasoning can be applied to each image by asking logical questions—for example, ’What if the key doesn’t open the door?’ The CTU unit proposes a new action:"the person is nervous" , then a new mental image is then generated. The same process is repeated and a new possible actions are generated: "the person breaks down the door", "the person call the firefighters for they open the door". Mental images generated are shown by figure 8.
@@ 7 Conclusion
In this paper, we addressed the ambitious challenge of enabling machine thinking by proposing a comprehensive framework inspired by recent advancements in neuroscience, deep learning, and artificial intelligence. While significant progress has been made in developing systems capable of processing sensory data, reasoning, and interacting with humans, our approach goes beyond existing models by integrating the concept of a Mental Image to simulate machine cognition.